OpenStudy (precal):

The equation x-(1/2)=-2(y+4) represents the equation of the tangent line to the graph of g(x) when x=-1. What is the value of g'(-1)?

OpenStudy (dumbcow):

put tangent line in "y=mx+b" form, then g'(x)= m since slope of tangent equals derivative

OpenStudy (dumbcow):

sorry g'(-1)

OpenStudy (precal):

ok let me do that

OpenStudy (precal):

I got -3 but it is not listed for my choices

OpenStudy (precal):

I have -1/2, 4, -2, and 1/2 and can not be determined

OpenStudy (dumbcow):

first did you divide by -2 ? --> y+4 = -x/2 +1/4

OpenStudy (precal):

yes and then I subtracted 4 to both sides

OpenStudy (precal):

I got y=(-1/2)x-(15/4)

OpenStudy (dumbcow):

good, now what is the slope?

OpenStudy (precal):

slope is -1/2 oh is that what they want, man I got to get better at reading these problems

OpenStudy (dumbcow):

haha yeah

OpenStudy (precal):

thanks you are the best